# Hydrogen reaction rate modeling based on convolutional neural network for large eddy simulation

### v1

@@ 1 Introduction
Hydrogen does not only challenge the way combustion chambers should be designed, it also constitutes a key difficulty for our simulation methods. Indeed, H2 is a very peculiar fuel due to its small Lewis number. Its high molecular diffusion, much faster than heat, gives rise to unique phenomena, especially when burned lean (i.e. with excess air) and premixed. First, lean hydrogen flames are prone to thermodiffusive instabilities. The disparity between heat and H2 mass fluxes amplifies small flame front perturbations, inducing significant flame wrinkling and strong variations of the reaction rates along the flame front [30, 3, 4]. Second, burning rates increase significantly with the strain rate. With H2 diffusing faster than heat, the reaction zone moves to the hot gas side of the flame as the strain increases. The reaction rates then increase monotonically with the strain rate, up to the extinction limit [28].
Unfortunately, these phenomena can occur on scales smaller than those that can be resolved numerically. The community nowadays uses Large Eddy Simulation (LES) methods, which involve simulating only the largest flow structures, to compute real complex configurations at an affordable computing cost. The smaller scales are modeled, according to Subgrid-Scale (SGS) models [30], also called subfilter-scale models. The complexity of these phenomena, specific to H2, poses many challenges to the development of turbulent combustion models for the simulation of premixed hydrogen flames that need to be tackled. Turbulent combustion can be approached via the “laminar flamelet concept” [27] in most practical cases. This concept assumes that a turbulent flame brush consists of a collection of laminar flames—called flamelets—wrinkled and stretched by the action of turbulent eddies. This hypothesis is verified when chemistry (as compared to transport processes) is fast such that it occurs in asymptotically thin layers. This has led to the development of so-called Flame Surface Density (FSD) models [31, 24, 7, 38], where the filtered burning rate $\overline{\dot{\omega}}$ is defined as the product of the flame surface density $\Sigma$ (i.e. the flame surface area per unit volume) by the local burning rate per unit of flame area:
@@
$\overline{\dot{\omega}}=\Sigma\left<\dot{\Omega}\right>_{s}\,\text{,}$ (1.1)
where $\dot{\Omega}\right>_{s}$ is the average burning rate along the flame surface. The modeling challenge is then decomposed into the problem of determining the unresolved surface area of the thin flame front, and the burning rate per unit area of that surface. H2, however, impacts on both these factors at the same time: instabilities generate flame surface at the SGS level, and preferential diffusion effects cause the burning rate to vary significantly along the flame front. Aniello et al. [1] have recently attempted to solve this issue by multiplying the burning rate by a corrective factor, determined from DNS [3, 4], decoupling the enhancement of the burning rate due to thermodiffusive instabilities and to usual turbulence wrinkling. However, this is no longer possible when the wrinkling scales induced by turbulent eddies become smaller than the characteristic lengths at which the instabilities occur. Moreover, Berger et al. [5] have reported a complex and synergistic interaction between thermodiffusive instabilities and turbulence, casting doubt on the possibility of decoupling these phenomena. A model capable of accounting for SGS thermodiffusive instabilities, turbulence wrinkling, stretching and their interactions in a general way for any configuration is needed for high-fidelity LES of H2 lean premixed flames.
@@
The phenomena involved being extremely complex, not well understood, and still subject to discussion, we propose to change the paradigm by developing a data-driven method based on Machine Learning (ML) instead of trying to model turbulent H2 burning rates by physical description. This work is motivated by the positive results of such an approach for the approximation of the subfilter flame surface density both a priori and a posteriori for methane-air combustion [22, 21]. The strategy is adapted to hydrogen combustion by training a Neural Network (NN) to estimate the H2 burning rate directly. In this way, no assumptions are made about the turbulent combustion regime, and the model can theoretically be applied to any combustion regime encountered during training. Firstly, a database of fully resolved turbulent H2-air premixed flames is generated using Direct Numerical Simulation (DNS), taking into account preferential and differential diffusion effects. The database includes five different equivalence ratios to assess the ability of the model to account for air/fuel ratio variation. The database is then filtered and downsampled to emulate data from LES. Three different filter sizes are used to assess the ability of the model to account for subfilter effects at different scales. From the emulated LES fields, a Convolutional Neural Network (CNN) is trained to estimate the LES burning rates based on known LES scalar quantities such as the filtered mixture fraction and progress variable. Once trained, the model is found to approximate the burning rates on solutions never seen during training with high accuracy. In addition to this, the model is able to approximate burning rates on filter sizes and equivalence ratios other than those used for training.
@@ 2.1 Problem statement
The simulation of turbulent flows at high Reynolds numbers cannot generally be performed with sufficient resolution to represent all the scales of the flow. LES aims to simulate only a part of the turbulent spectrum to achieve a reasonable computational cost. Scales are cut by spatial filtering, and unresolved scales are modeled using so-called subfilter-scale models. In the LES context, the Favre-filtered transport equation for the mass fraction of the $k^{\text{th}}$ species $Y_{k}$ reads
$\frac{\partial\overline{\rho}\widetilde{Y_{k}}}{\partial t}+\frac{\partial}{\partial x_{i}}\left(\overline{\rho}\widetilde{Y_{k}}\widetilde{u_{i}}\right)=-\frac{\partial}{\partial x_{i}}\left[\overline{J_{i,k}}+\overline{J_{i,k}}^{t}\right]+\overline{\dot{\omega}_{k}}\;\mathrm{,}$ (2.1)
where $\rho$ is the gas density, $u_{i}$ is the $i^{\mathrm{th}}$ velocity component, $\overline{J_{i,k}}$ is the filtered diffusive species flux, $\overline{J_{i,k}}^{t}$ is the subfilter species flux, and $\overline{\dot{\omega}_{k}}$ the filtered chemical source term of the $k^{\mathrm{th}}$ species. $\overline{\varphi}$ and $\widetilde{\varphi}$ denotes spatial and Favre (i.e. mass weighted) filtering of $\varphi$ with $\widetilde{\varphi}=\overline{\rho\varphi}/\overline{\rho}$.
Modeling filtered chemical source terms has challenged the scientific community for many years. Although efforts are considerable, there is currently no universal method that covers all fuels and all combustion regimes, given the complexity of the associated physics. The laminar flamelet concept [27] is the prevalent framework for the derivation of premixed turbulent combustion models. In most practical applications, combustion takes place in thin layers embedded within the turbulent flow field and this concept is valid. The lean premixed combustion of H2 nevertheless challenges the modeling of the burning rates along the flame front at a subfilter-scale level. As discussed in the introduction, the high mobility of H2 causes pronounced fluctuations in reaction rates along the flame front. A precise description of the preferential and differential diffusion, turbulence wrinkling, and their interaction [5] at subfilter-scale level is missing.
In this work, we evaluate the possibility of training a ML-based model to approximate the filtered reaction rates as a function of other LES scalar quantities. The basic ingredients to describe turbulent flames remain the quantities introduced for laminar flame analysis: the progress variable $c$ and the mixture fraction $\xi$. The progress variable equals zero in fresh gas, one in burnt gas, and varies monotonically during combustion. Many LES modeling methods replace Eq. (2.1) with an equation for $\widetilde{c}$ (e.g. Ref. [7]). However, solving a transport equation for a filtered progress variable involves additional complications to account for the very strong preferential and differential diffusion effects present in $\mathrm{H_{2}}$-air mixtures [32]. Moreover, it has been shown that for lean hydrogen premixed flames, reaction rates cannot be correctly approximated solely as a function of the progress variable and must be approximated by a higher-dimensional manifold [19, 20]. The fast molecular diffusion of H2 with respect to O2 leads to locally leaner and richer mixtures along the flame front and hence, induces strong reaction rate fluctuations. An additional parameter such as the mixture fraction $\xi$ or the equivalence ratio $\phi$ is required to capture local variations in mixture composition. We then propose to solve Eq. (2.1) for four major species: $\mathrm{H_{2}}$, $\mathrm{O_{2}}$, $\mathrm{H_{2}O}$ and $\mathrm{N_{2}}$. The transport of these species makes it possible to account for the rapid diffusion of $\mathrm{H_{2}}$, to estimate a local mixture fraction $\xi$, and to define a progress variable $c$ based on the species. The filtered mixture fraction $\widetilde{\xi}$ is based on Bilger’s definition [6]
$\widetilde{\xi}=\frac{\widetilde{\beta}-\beta^{o}}{\beta^{f}-\beta^{o}}\;\text{with}\;\beta=\frac{Z_{\mathrm{H}}}{2W_{\mathrm{H}}}-\frac{Z_{\mathrm{O}}}{W_{\mathrm{O}}}\,\text{,}$ (2.2)
@@
where $Z_{m}$ and $W_{m}$ are the elemental mass fraction and atomic mass of the element $m$. The superscripts $o$ and $f$ denote oxidizer and fuel stream. The filtered progress variable is defined as
@@
[…]
$\overline{\dot{\omega}_{\mathrm{N_{2}}}}=0\;\text{and}\;\overline{\dot{\omega}}=-\overline{\dot{\omega}_{\mathrm{H_{2}}}}=-2\frac{W_{\mathrm{H_{2}}}}{W_{\mathrm{O_{2}}}}\overline{\dot{\omega}_{\mathrm{O_{2}}}}=\frac{W_{\mathrm{H_{2}}}}{W_{\mathrm{H_{2}O}}}\overline{\dot{\omega}_{\mathrm{H_{2}O}}}\,\text{,}$ (2.4)
where $W_{k}$ is the molecular weight of the species $k$. In this framework, the challenge is now to model $\overline{\dot{\omega}}$ as a function of $\widetilde{c}$ and $\widetilde{\xi}$ fields.
@@ 2.3 Data generation and training strategy
    From the filtered variables, $\widetilde{c}$, $\widetilde{\phi}$ and $\overline{\dot{\omega}}$ (Eqs (2.3), (2.6) and (2.4)) are calculated.
@@
    It is important that the CNN does not learn the specific configuration of the slot burner, but only learns generic flame elements. Therefore, training is not carried out on the entire domain, but rather on randomly selected cubes in the domain as in Refs [22, 21]. A number of cubes per solution $N_{c}=40$ is randomly extracted for training and validation. The size of the cubes is $16\times 16\times 16$ points.
@@ 2.3.1 Direct numerical simulation
DNS of the slot burner cases (Table 1) are performed using the AVBP [35, 13] massively parallel code solving the compressible multi-species Navier-Stokes equations. A third order accurate Taylor–Galerkin scheme is adopted for discretization of the convective terms [9]. Navier-Stokes Characteristic Boundary Conditions [29] are imposed at the inlets and at the outlet. Viscosity follows a power law function of temperature. Thermal diffusivity is computed from the viscosity using a constant Prandtl number: $\mathrm{Pr}=0.66$. Species diffusion fluxes rely on Hirschfelder and Curtis approximation [16] where diffusion coefficients are computed using a constant Schmidt number specific for each species (Table 2). This approach takes into account non-unity Lewis number and differential diffusion between the different species. Hydrogen chemical kinetics relies on the San Diego mechanism [34], already successfully used for H2-air premixed combustion in Ref. [10]. This mechanism comprises $9$ species and $21$ reactions.
@@
The mesh is a homogeneous cartesian grid with constant element size $\Delta_{x}=0.1$ mm. This gives a resolution index $\mathrm{RI}=\delta_{th,L}^{0}/\Delta_{x}+1$ that varies between $12.5$ and $5.2$ for $\phi_{g}$ between $0.35$ and $0.7$. The length of the domain in the $x$-direction $L_{x}$ is adapted to the length of turbulent the flame brush. It varies from $76$ mm for $\phi_{g}=0.35$ to $36$ mm for $\phi_{g}=0.7$ (Fig. 2). The number of grid points accordingly varies from $39$ to $19$ million. Calculations were carried out on the new Alps Research Infrastructure at the Swiss National Supercomputing Centre (CSCS) using AMD EPYC 7742 CPUs. A total of $1.6$ million core-hours were used to generate the database. Depending on the grid size, computations are performed using between $5,120$ and $2,560$ computing cores with good efficiency thanks to the massively parallel computing ability of the AVBP code.

### v2

@@ 1 Introduction
The community nowadays uses Large Eddy Simulation (LES) methods, which involve simulating only the largest flow structures, to compute real complex configurations at an affordable computing cost. The smaller scales are modeled, according to subfilter-scale models [42, 19]. In LES, a filtered quantity $\varphi$ is defined as
[…]
where $F$ is the LES filter. Balance equations are obtained by applying the filter $F$ to the multispecies Navier-Stokes equations. The filtered transport equation of a reactive scalar $\varphi$ reads
$\partial_{t}\left(\overline{\rho}\widetilde{\varphi}\right)+\nabla\cdot\left(\overline{\rho}\widetilde{\boldsymbol{u}}\widetilde{\varphi}\right)-\nabla\cdot(\overline{\rho}\widetilde{D}_{\varphi}\nabla\widetilde{\varphi})=\nabla\cdot\Gamma_{\varphi}+\overline{\dot{\omega}}_{\varphi}\text{ ,}$ (1.2)
where $\widetilde{\cdot}$ denotes density-weighted Favre filtering (i.e. ~$\widetilde{\cdot}=\overline{\rho\cdot}/\overline{\rho}$), with density $\rho$, time $t$, velocity $\boldsymbol{u}$. $D_{\varphi}$, $\Gamma_{\varphi}$ and ${\dot{\omega}}_{\varphi}$ are the molecular diffusivity, subfilter-scale flux and source term of $\varphi$. The filtering operation produces unclosed terms located on the RHS of Eq. (1.2) that must be modeled. The subfilter-scale flux comprises the transport of $\varphi$ by unresolved fluctuations of molecular diffusive flux and momentum
$\Gamma_{\varphi}=\Gamma_{\varphi,D}-\Gamma_{\varphi,C}\;\;\text{with}\;\Gamma_{\varphi,D}=\overline{\rho D_{\varphi}\nabla\varphi}-\overline{\rho}\widetilde{D}_{\varphi}\nabla\widetilde{\varphi}\;\;\text{and}\;\Gamma_{\varphi,C}=\overline{\rho\mathbf{u}\varphi}-\overline{\rho}\widetilde{\mathbf{u}}\widetilde{\varphi}\;\text{.}$ (1.3)
It is usually modeled neglecting $\Gamma_{\varphi,D}$ and involving gradient transport models for $\Gamma_{\varphi,C}$ with a turbulent viscosity hypothesis [11]. Countergradient phenomenon can be taken into account via more complex modeling [53, 52], although a large part will be described at the resolved scale in LES [11]. Alternatively, $\Gamma_{\varphi}$ can be tabulated within the filtered tabulated chemistry framework [20] or modeled via Machine Learning (ML) algorithms [49]. Modeling of the filtered chemical source term $\overline{{\dot{\omega}}}_{\varphi}$ is extremely challenging because combustion generally occurs within reaction fronts that are smaller than the LES computational mesh. Furthermore, there is a strong interaction between turbulence and chemistry: chemical heat release influences turbulence, while turbulence, in turn, modifies flame structure, which can either enhance or inhibit chemical reactions.
To overcome these difficulties, several approaches have been proposed in the context of premixed combustion, which is the focus of this paper. A common technique is to artificially thicken the reaction front, enabling it to be resolved on the given computational mesh. This is the basis of the Thickened Flame (TF) model, which achieves this by increasing all diffusivities by a factor $F$ while decreasing all reaction rates by the same factor to maintain the correct flame propagation speeds [12]. The LES modeling then focuses on an efficiency function $E$ taking into account the effects of thickening on flame propagation, such as the loss of flame wrinkling [14, 13, 42]. Another approach consists of filtering the flame front with a LES filter larger than the mesh size [11, 17]. Filtering is usually done on the equation of a reaction progress variable $c$ representing the evolution of combustion within the flame front. The balance equation for $\widetilde{c}$ may be recast as a level set G-equation [27]
@@
$\partial_{t}\left(\bar{\rho}\widetilde{c}\right)+\nabla\cdot\left(\bar{\rho}\widetilde{\boldsymbol{u}}\widetilde{c}\right)+\nabla\cdot\left(\overline{\rho\boldsymbol{u}c}-\overline{\rho}\widetilde{\boldsymbol{u}}\widetilde{c}\right)=\nabla(\overline{\rho D_{c}\nabla\cdot c})+\overline{\dot{\omega}}_{c}=\overline{\rho s_{d}|\nabla c|}\;\text{,}$ (1.4)
which gives rise to a flame front displacement term at speed $s_{d}$, $\overline{\rho s_{d}|\nabla c|}$, which requires modeling. Another way to close the equation is to tabulate all the closure terms from filtered laminar premixed flames [20], assuming the turbulent flame can be represented as an ensemble of locally One-Dimensional (1D) laminar flamelets [38]. In all cases, a model to retrieve the subfilter-scale effects on the thickened/filtered flame front is needed.
Modeling subfilter-scale phenomena is particularly challenging for hydrogen combustion. Indeed, H2 is a very peculiar fuel due to its small Lewis number. Its high molecular diffusion, much faster than heat, gives rise to unique phenomena, especially when burned lean (i.e. with excess air) and premixed. First, lean H2 flames are prone to thermodiffusive instabilities. The disparity between heat and H2 mass fluxes (called differential diffusion) amplifies small flame front perturbations, inducing significant flame wrinkling and strong variations of the reaction rates along the flame front [42, 5, 6]. Second, burning rates increase significantly when lean H2 flame is subjected to a strain [39, 54], up to a certain limit. The resulting increase in flame speed can be substantial. For example, the experimental work in Ref. [54] reports a four-fold enhancement in flame speed compared to an unstrained flame.
These phenomena invalidate usual Flame Surface Density (FSD) approaches [43, 34, 11, 51] based on the assumption that subfilter-scale effects are limited to an increase in flame surface density through wrinkling due to turbulence alone. Aniello et al. [1] have recently attempted to address this issue by multiplying the filtered chemical source terms by a corrective factor, determined from Direct Numerical Simulation (DNS) [5, 6], decoupling the enhancement of combustion rates due to thermodiffusive instabilities and to usual turbulence wrinkling. However, this is no longer possible when the wrinkling scales induced by turbulent eddies become smaller than the characteristic lengths at which the instabilities occur. Moreover, Berger et al. [7] have reported complex and synergistic interactions between thermodiffusive instabilities and turbulence, casting doubt on the possibility of decoupling these phenomena. Gaucherand et al. [22] attempted to correct for unresolved straining effects by multiplying the filtered chemical source terms by a strain-to-unstrained flame speed ratio determined using 1D flames. The strain rate was assumed to be constant over the entire domain at a value of $2000$ $\mathrm{s^{-1}}$, not taking into account the very wide strain variation along the turbulent flame. Additional efforts must be made to account for local strain effects, which is extremely difficult in LES because the information on the strain that a turbulent flame actually undergoes at the subfilter-scale level is unknown. Subfilter Probability Density Function (PDF) modeling for H2 turbulent combustion is also being developed [8, 40, 57, 18], where the subfilter-scale distribution of thermochemical variables is usually modeled using $\beta$-PDF. The turbulence-chemistry interactions are described through the subfilter PDF while the flamelet assumption [38] links local thermochemical variables to those of the laminar flame elements. Thus, a set of 1D laminar flames is needed to be computed, assuming a specific flame archetype a priori. Defining the subfilter PDF is a particular challenge. For example, Refs [8, 57] use a Dirac function for the mixture fraction distribution, while it is reported in Ref. [40] that it leads to modeling errors. Indeed, preferential diffusion in H2-air mixtures can lead to fluctuation of the mixture fraction at the subfilter-scale level. Furthermore, the effect of unstable modes that occur partially at the subfilter-scale must be included in the subfilter PDF to avoid underestimating the flame consumption speed [30, 40].
@@
The phenomena involved in H2 turbulent combustion being extremely complex, not well understood, and still subject to discussion, we propose to change the paradigm by developing a data-driven method based on ML instead of trying to model turbulent H2 burning rates by physical description. This work is motivated by the positive results of such an approach for the approximation of the subfilter flame surface density both a priori and a posteriori for methane-air combustion [32, 31, 26]. The strategy is adapted to hydrogen combustion by training a Neural Network (NN) to estimate the H2 burning rate directly. In this way, no assumptions are made about the turbulent combustion regime, and the model can theoretically be applied to any combustion regime encountered during training. Firstly, a database of fully resolved turbulent H2-air premixed flames is generated using DNS, taking into account preferential and differential diffusion effects. The database includes five different equivalence ratios to assess the ability of the model to account for air/fuel ratio variation. The database is then filtered and downsampled to emulate data from LES. Three different filter sizes are used to assess the ability of the model to account for subfilter effects at different scales. From the emulated LES fields, a Convolutional Neural Network (CNN) is trained to estimate the LES burning rates based on known LES scalar quantities such as the filtered mixture fraction and progress variable. Once trained, the model is found to approximate the burning rates on solutions never seen during training with high accuracy. In addition to this, the model is able to approximate burning rates on filter sizes and equivalence ratios other than those used for training.
@@ 2.1 Problem statement
The simulation of turbulent flows at high Reynolds numbers cannot generally be performed with sufficient resolution to represent all the scales of the flow. LES aims to simulate only a part of the turbulent spectrum to achieve a reasonable computational cost. Scales are cut by spatial filtering, and unresolved scales are modeled using so-called subfilter-scale models. In the LES framework, the Favre-filtered transport equation for the mass fraction of the $k^{\text{th}}$ species $Y_{k}$ follows Eq. (1.2). As discussed in the introduction, modeling the filtered chemical source terms in LES of reacting flows is extremely challenging because chemistry largely occurs at the subfilter-scale level in most cases. This is further complicated in lean premixed H2-air mixtures because of thermodiffusive effects, causing additional flame wrinkling and pronounced fluctuations in reaction rates along the flame front. A precise description of the preferential and differential diffusion, turbulence wrinkling, and their interactions [7] at subfilter-scale level is required.
In this work, we focus on the development of a data-driven model for the filtered chemical source term of H2 in lean premixed turbulent combustion. ML is used to discover structural relations between resolved scalar quantities and subfilter-scale turbulence-chemistry interactions. We develop this method for the H2 filtered “burning rate” $\overline{\dot{\omega}}$, defined as
$\overline{\dot{\omega}}=-\overline{\dot{\omega}}_{\mathrm{H_{2}}}\text{ ,}$ (2.1)
with the aim of establishing an ML-based modeling framework and a proof of concept for H2 turbulent combustion. Final implementation will depend on the given LES modeling framework. For example, the ML model could replace the tabulation method proposed in Refs [29, 30] for the chemical source term of the progress variable. This is what has been done in Ref. [49] with direct mapping of the closure terms for a methane-air flame. The ML framework could also be based on a database with one-step chemistry [35, 46] to close the LES multi-species transport equations.
The basic ingredients to describe turbulent flames remain the quantities introduced for laminar flame analysis: the progress variable $c$ and the mixture fraction $\xi$. The progress variable equals zero in fresh gas, one in burnt gas, and varies monotonically during combustion. Mixture fraction $\xi$ or the equivalence ratio $\phi$ is used to inform the ML-based model of the composition of the mixture, which impacts burning rates. The mixture fraction is based on Bilger’s definition [9]
${\xi}=\frac{{\beta}-\beta^{o}}{\beta^{f}-\beta^{o}}\;\text{with}\;\beta=\frac{Z_{\mathrm{H}}}{2W_{\mathrm{H}}}-\frac{Z_{\mathrm{O}}}{W_{\mathrm{O}}}\,\text{,}$ (2.2)
@@
where $Z_{m}$ and $W_{m}$ are the elemental mass fraction and atomic mass of the element $m$. The superscripts $o$ and $f$ denote oxidizer and fuel stream. The filtered progress variable $\widetilde{c}$ is defined as
@@
where superscripts $u$ and $b$ denote values in unburnt and burnt gas and $\xi_{s}$ is the mixture fraction value at stoichiometry. We have arbitrarily chosen to base the progress variable on the H2 mass fraction. It is shown in supplementary material that it is also possible to base the progress variable on the H2O mass fraction without any impact on the training and prediction accuracy of the model. In this framework, the challenge is now to model $\overline{\dot{\omega}}$ as a function of $\widetilde{c}$ and $\widetilde{\xi}$ fields.
@@ 2.3 Data generation and training strategy
    The filtered progress variable $\widetilde{c}$ is calculated using Eq. (2.3) with $\widetilde{Y}_{\mathrm{H_{2}}}=\overline{\rho Y_{\mathrm{H_{2}}}}/\overline{\rho}$ and $\widetilde{\xi}=\overline{\rho\xi}/\overline{\rho}$. The filtered equivalence ratio $\widetilde{\phi}$ is calculated using Eq. (2.5) with $\widetilde{\xi}=\overline{\rho\xi}/\overline{\rho}$. The filtered burning rate $\overline{\dot{\omega}}$ is calculated using Eq. (2.1) with $\overline{\dot{\omega}}_{\mathrm{H_{2}}}$.
@@
    It is important that the CNN does not learn the specific configuration of the slot burner, but only learns generic flame elements. Depending on the effective receptive field of the CNN, training using the emulated LES solutions directly could encode in the network information on the distance to the inlets, outlet or the wall for example. This would drastically reduce the CNN ability to generalize, and must be avoided. To address this, training is performed on randomly selected cubes in the domain as in Refs [32, 31]. This also enables the CNN to be invariant to translation [10].
@@ 2.3.1 Direct numerical simulation
DNS of the slot burner cases (Table 1) are performed using the AVBP [48, 23] massively parallel code solving the compressible multi-species Navier-Stokes equations. A third order accurate Taylor–Galerkin scheme is adopted for discretization of the convective terms [15]. Navier-Stokes Characteristic Boundary Conditions [41] are imposed at the inlets (relaxation factor of $1000$ $\mathrm{s^{-1}}$) and at the outlet (relaxation factor of $200$ $\mathrm{s^{-1}}$). Dynamic viscosity $\mu$ follows a power law function of temperature $T$
$\mu=\mu_{0}\left(\frac{T}{T_{0}}\right)^{\gamma}\text{ ,}$ (2.7)
with $\mu_{0}=8.062\times 10^{-5}$ $\mathrm{kg/m/s}$, $T_{0}=2.645\times 10^{3}$ K and $\gamma=6.481\times 10^{-1}$. Thermal diffusivity is computed from the viscosity using a constant Prandtl number: $\mathrm{Pr}=0.66$. Species diffusivities are computed using a constant Schmidt number specific for each species (Table 2). This approach takes into account non-unity Lewis numbers and preferential diffusion between the different species. It is verified that the errors made by the simplified transport description are negligible by comparing the results with simulations using a mixture-averaged transport model, see supplementary material. Soret and Dufour transport processes are ignored in the simulations of the present work. Including these transport processes could enhance thermodiffusive instabilities, increase source terms and flame surface curvature [47]. Although this would require another training on the new database, it does not affect the conclusions of this work. Indeed, it has been shown that neglecting these effects still leads to simulations that capture much of the peculiar behaviour of lean premixed H2-air flames including thermodiffusive instabilities and enhanced burning rates [47, 4, 3, 40]. Hydrogen chemical kinetics relies on the San Diego mechanism [45], already successfully used for H2-air premixed combustion in Ref. [16]. This mechanism comprises $9$ species and $21$ reactions.
@@
The mesh is a homogeneous cartesian grid with constant element size $\Delta_{x}=0.1$ mm. This gives a resolution index $\mathrm{RI}=\delta_{th,L}^{0}/\Delta_{x}+1$ that varies between $12.5$ and $5.2$ for $\phi_{g}$ between $0.35$ and $0.7$. $\delta_{th,L}^{0}$ is the laminar thermal flame thickness. A mesh sensitivity study was carried out where we verified that turbulent combustion statistics are not affected by mesh refinements. This study is included as supplementary material. The length of the domain in the $x$-direction $L_{x}$ is adapted to the length of turbulent the flame brush. It varies from $76$ mm for $\phi_{g}=0.35$ to $36$ mm for $\phi_{g}=0.7$ (Fig. 2). The number of grid points accordingly varies from $39$ to $19$ million. Calculations were carried out on the new Alps Research Infrastructure at the Swiss National Supercomputing Centre (CSCS) using AMD EPYC 7742 CPUs. A total of $1.6$ million core-hours were used to generate the database. Depending on the grid size, computations are performed using between $5,120$ and $2,560$ computing cores with good efficiency thanks to the massively parallel computing ability of the AVBP code.